OpenStudy (anonymous):

Rearrange this equation to isolate c a=b(1/c - 1/d)

OpenStudy (anonymous):

divide by b

OpenStudy (anonymous):

so itd be...?

OpenStudy (anonymous):

excuse me, divide BOTH sides by b

OpenStudy (anonymous):

I've got this so far: a/b +1/d= 1/c

OpenStudy (anonymous):

I can't remember how to get past here

OpenStudy (anonymous):

alright so

OpenStudy (anonymous):

multiply both sides by c c(a/b+1/d)=1/c (c)

OpenStudy (anonymous):

I considered that but then I got confused

OpenStudy (anonymous):